OpenStudy (anonymous):

write a quadratic equation having the given solution -2, 8 please help

OpenStudy (ajprincess):

solutions given So x=-2 or x=8 (x+2)(x-8)=0 distribute this and find the quadratic equation. does that help?

OpenStudy (anonymous):

so that mean I have to solve for x

OpenStudy (ajprincess):

(x+2)(x-8)=0 x(x-8)+2(x-8)=0 distribute and simplify

OpenStudy (anonymous):

how do I distribute and simplify it please i'm sorry i am really bad with math

OpenStudy (ajprincess):

x(x-8)+2(x-8)=0 x*x+x*8+2*x-2*8=0 x^2+8x+2x-16=0 Combine the like terms

OpenStudy (anonymous):

thank you